A house was purchased for $200,000 with a loan equal to 85% of the purchase price and the price is expected to rise to $205,000 at the end of the year. What is the Expected Appreciation rate on Home Equity (rounded)?

10%

2.5%

17%

8%
Answer #1
Loan amount = Purchase price * Loan % = 200000 * 85% 170000
Home equity = Purchase price - Loan amount = 200000 - 170000 30000
Expected Appreciation amount = 205000 - 200000 5000
Expected appreciation rate = Expected appreciation amount / Home equity = 5000 / 30000 17%